OpenStudy (anonymous):

a number cube with faces labeled 1-6 rolled 72 times.about how many times would you expect the result to be 4

OpenStudy (anonymous):

The probability of rolling a 4 is 1/6th. \[P(4)=\frac{1}{6}\] so we would expect that a 4 would come up about 1 out of every 6 rolls. For 72 rolls we would expect to see \[\frac{72}{6}=12\] rolls with a 4.